### v1

# Reentrant Superconductivity in Zeeman Fields

Abstract
We propose a theoretical model for a superconductor that exhibits the reentrant superconductivity in Zeeman fields. The Bogoliubov-de Gennes Hamiltonian includes three vectors in spin space: a $d$ vector of a spin-triplet superconducting state, a potential representing spin-orbit interactions, and a Zeeman field. When the three vectors are perpendicular to one another, the spin-orbit interaction suppress superconductivity in weak Zeeman fields and enhances superconductivity in strong Zeeman fields. The instability (stability) of superconducting state is characterized by the appearance of odd-frequency (even-frequency) Cooper pairs.
Introduction: An external magnetic field suppresses superconductivity in two ways. The orbital effect fluctuates the phase of superconducting condensate and increases the free-energy of a superconducting state. This effect is common for all superconductors (SCs). In layered SCs or monolayer SCs, the orbital effect is negligible for a magnetic field applied parallel to the layers. Only the Zeeman effect modifies the superconducting states. A Zeeman field always suppresses spin-singlet superconductivity. The critical magnetic field $H_{c}$ at zero temperature is called Pauli limit $H_{\mathrm{P}}$[14, 4]. For spin-triplet superconductivity, the effect depends on the alignment between the direction of a Zeeman filed $\bm{H}$ and that of a $\bm{d}$ vector. A Zeeman filed parallel to $\bm{d}$ suppresses superconductivity, whereas that perpendicular to $\bm{d}$ does not affect the superconducting state. In experiments, Zeeman-field-induced superconductivity (ZFIS) or reentrant superconductivity have been observed various materials such as an organic SC $\lambda-\mathrm{(BETS)}_{2}\mathrm{FeCl}_{4}$ [20], Europium compounds [10], UCoGe[2], and UTe2 [12, 1, 7]. The last one has attracted attention these days as a promising candidate of a spin-triplet SC. However, our understanding of ZFIS remains extremely limited. Only the Jaccarino-Peter compensation [8] is the widely accepted as the mechanism explaining the transition from a magnetic phase to a superconducting phase. Thus, another scenario for explaining the phenomenon is desired to understand physics behind the experimental findings. We will address such an issue in this paper.
To demonstrate the ZFIS, we introduce the third vector in spin space $\bm{\alpha}$ that describes spin-orbit interactions (SOIs). A Zeeman field and a SOI induce additional pairing correlations. Although they are not linked to the pair potentials, they govern the stability of superconductivity depending on their frequency symmetry classes. Odd-frequency Cooper pairs decrease the transition temperature $T_{c}$ [3, 11, 19] because they decrease the superfluid weight. On the other hand, even-frequency Cooper pairs increase $T_{c}$[16] In this paper, we show that spin-singlet Cooper pairs whose correlation function is proportional to a scaler chiral product $\bm{\alpha}\times\bm{H}\cdot\bm{d}$ assist spin-triplet superconductivity in sufficiently strong Zeeman fields. We use the unit of $k_{B}=\hbar=c=1$, where $k_{B}$ is the Boltzmann constant and $c$ is the speed of light.
Figure 1: The critical magnetic field $H_{c}$ is plotted as a function of temperature $T$. (a) All the three vectors are align in parallel, $\bm{d}\parallel\bm{\alpha}\parallel\bm{H}$. The arrow at the vertical axis indicates the Pauli limit $H_{\mathrm{P}}$. (b) The three vectors are perpendicular to one another, $\bm{d}\perp\bm{\alpha}$, $\bm{d}\perp\bm{H}$, and $\bm{\alpha}\perp\bm{H}$. At $H=0$, superconductivity is absent because of $\alpha>\mu_{\mathrm{B}}H_{\mathrm{P}}$. (c) The mixed states with $\alpha_{x}=0.75T_{0}$ exhibit the reentrant superconductivity. The two superconducting phases tend to separate from each other for large $\alpha_{y}$.
Model: We consider a electronic structure in two dimension, where two Fermi surfaces surrond $\bm{K}$ and $\bm{K}$ in momentum space. We describe such two Fermi surfaces in terms of two valleys. The normal state Hamiltonian is given by
$\check{H}_{\mathrm{N}}$ $(\bm{k})=\hat{\xi}_{\bm{k}}\hat{\sigma}_{0}+\bm{\alpha}\cdot\hat{\bm{\sigma}}\hat{\rho}_{z}+\mu_{\mathrm{B}}\bm{H}\cdot\hat{\bm{\sigma}}\hat{\rho}_{0},$ (1)
$\hat{\xi}_{\bm{k}}$ $=\frac{1}{2m}\left(\bm{k}\hat{\rho}_{0}-\bm{K}\hat{\rho}_{z}\right)^{2}-\mu,$ (2)
where $\mu$ is the chemical potential and $\bm{H}$ represents a Zeeman field with $\mu_{\mathrm{B}}$ being the Bohr magneton. The Pauli matrices in spin and valley spaces are denoted by $\hat{\bm{\sigma}}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$ and $\hat{\bm{\rho}}=(\hat{\rho}_{x},\hat{\rho}_{y},\hat{\rho}_{z})$, respectively. The unit matrices in the two spaces are denoted by $\hat{\sigma}_{0}$ and $\hat{\rho}_{0}$. We assume that the SOI $\bm{\alpha}$ is independent of $\bm{k}$ and changes its sign in the two valleys. The time-reversal operation is given by $\mathcal{T}=i\hat{\sigma}_{y}\hat{\rho}_{x}\mathcal{K}$, where $\mathcal{K}$ means the complex conjugation plus $\bm{k}\to-\bm{k}$. The particle-hole conjugation is represented by $\undertilde{\check{H}}_{\mathrm{N}}(\bm{k})=\check{H}^{*}_{\mathrm{N}}(-\bm{k})$ as usual. Two electrons at the different valleys form a spin-triplet Cooper pair. The pair potential for such a pair is represented as
$\check{\Delta}=i\,\bm{d}\cdot\hat{\bm{\sigma}}\,\hat{\sigma}_{y}\,i\hat{\rho}_{y},$ (3)
where $\bm{d}$ is independent of $\bm{k}$. Instead of $\bm{d}$ being odd-parity function, making valley-parity odd preserves the antisymmetric property of the pair potential derived from the Fermi-Dirac statistic of electrons. We note that all Cooper pairs in this model belong to even-parity $s$-wave symmetry class. Although the similar electronic structures are realized in transition-metal dicalcogenides [15, 21, 5], we do not focus on any specific materials in this paper. The advantage of the model is briefly summarized as follows. In single-band SCs, $\bm{\alpha}$ and $\bm{d}$ are an odd-parity function depends on $\bm{k}$. In our model, they are described by the $\bm{k}$ independent potentials which change sign under interchanging the valley indices. As a result, magnetic properties of a SC is purely derived from the relative vector configuration among $\bm{d}$, $\bm{\alpha}$, and $\bm{H}$. At the end of this paper, we will confirm that main conclusions of this paper are valid also in single-band odd-parity spin-triplet SCs.
We solve the Gor’kov equation for the Bogoliubov-de Gennes (BdG) Hamiltonian
$\left[i\omega_{n}-H_{\mathrm{BdG}}(\bm{k})\right]\begin{bmatrix}\check{\mathcal{G}}&\check{\mathcal{F}}\\ -\undertilde{\check{\mathcal{F}}}&-\undertilde{\check{\mathcal{G}}}\end{bmatrix}_{(\bm{k},\omega_{n})}=1,$ (4)
$H_{\mathrm{BdG}}(\bm{k})=\begin{bmatrix}\check{H}_{\mathrm{N}}(\bm{k})&\check{\Delta}\\ -\undertilde{\check{\Delta}}&-\undertilde{\check{H}}_{\mathrm{N}}(\bm{k})\end{bmatrix}$, (5)
where $\omega_{n}=(2n+1)\pi T$ is a Matsubara frequency with $T$ being a temperature. As shown in Eq. (17) in Appendix A, the Hamiltonian is block-diagonalized into two particle-hole spaces. The analytical expression of the anomalous Green’s function is supplied in Eq. (19).
Table 1: Symmetry classification of the pair correlation functions. The top row corresponds to the spin-triplet pairing correlation that is linked to the pair potential through the gap equation. It appears at the first term in Eqs. (6) and (9). The second and third low represent odd-frequency pairing correlations appearing at the second term in Eqs. (6) and (9). They make the superconducting unstable. The cooperative effect between a Zeeman filed and a SOI induces the spin-singlet pairing correlation at the bottom row. All the Cooper pairs belong to even-momentum-parity $s$-wave symmetry class.
frequency spin ($\times i\hat{\sigma}_{y}$) valley-parity
even triplet $\bm{d}\cdot\hat{\bm{\sigma}}$ odd $\hat{\rho}_{y}$ pair potential
odd singlet $\bm{d}\cdot\bm{H}$ odd $\hat{\rho}_{y}$ $F_{\parallel}$
odd triplet $\bm{\alpha}\times\bm{d}\cdot\hat{\bm{\sigma}}$ even $\hat{\rho}_{x}$ $F_{\perp}$
even singlet $\bm{\alpha}\times\bm{d}\cdot\bm{H}$ even $\hat{\rho}_{x}$ $F_{\perp}$
Results: At first, we discuss a configuration $\bm{\alpha}\parallel\bm{H}\parallel\bm{d}$, where the three vectors align in the same direction. The anomalous Green’s function is given in Eqs (26). To solve the linearized gap equation, the summation over $\bm{k}$ is necessary. The shift of the wavenumber $\bm{k}\to\bm{k}\mp\bm{K}$ for $\xi_{\pm}$ remains the results unchanged because $\bm{\alpha}$ and $\bm{d}$ are independent of $\bm{k}$. By changing the summation over $\bm{k}$ to the integration over $\xi_{\bm{k}}\mp\alpha\hat{\rho}_{z}$, we reach an expression of
$\frac{1}{V_{\mathrm{vol}}}\sum_{\bm{k}}$ $\check{F}_{\parallel}(\bm{k},\omega_{n})=-\frac{N_{0}\pi}{|\omega_{n}|}\left[\frac{\omega_{n}^{2}}{\omega_{n}^{2}+\mu_{\mathrm{B}}^{2}H^{2}}\bm{d}\cdot\hat{\bm{\sigma}}\right$.
$\left.-\frac{i\,\omega_{n}\,\mu_{\mathrm{B}}\bm{d}\cdot\bm{H}}{\omega_{n}^{2}+\mu_{\mathrm{B}}^{2}H^{2}}\right]\left(i\hat{\sigma}_{y}\right)(i\hat{\rho}_{y}),$ (6)
where $N_{0}$ is the density of states at the Fermi level per spin. The SOI modifies the band-dispersion only slightly in this configuration. The first term contributes to the pair potential through the gap equation in Eq. (23). A Zeeman field decreases its amplitude and generates odd-frequency pairing correlation appearing at the second term. Odd-frequency pairing correlations makes the superconducting state unstable [3]. The two pairing correlations are listed at the first and second rows in Table 1. We find that the structure of the anomalous Green’s function in Eq. (6) is essentially the same as that for a spin-singlet superconductor under a Zeeman field by changing $\bm{d}\cdot\hat{\bm{\sigma}}\to\Delta$ and $\bm{d}\cdot\bm{H}\to\Delta\bm{H}\cdot\hat{\bm{\sigma}}$. In fact, the resulting linearized gap equations in the two cases
$\ln\frac{T}{T_{0}}+2\pi T\sum_{\omega_{n}>0}\frac{1}{\omega_{n}}\left[1-\frac{\omega_{n}^{2}}{\omega_{n}^{2}+\mu_{\mathrm{B}}^{2}H^{2}}\right]=0,$ (7)
are identical to each other. Here $T_{0}$ is the transition temperature at $\bm{H}=\bm{\alpha}=0$. We draw the HTH-T phase diagram shown in Fig. 1(a) in two ways: solving the gap equation and finding the minimum of the free-energy,
$\Omega_{S}=$ $-\frac{1}{2}T\frac{1}{V_{\mathrm{vol}}}\sum_{\bm{k}}\sum_{j=1}^{4}\log\left(2\cosh\left[\frac{E_{\bm{k},j}}{2T}\right]\right)+\frac{\bm{d}^{2}}{g}$
$+\text{Const.},$ (8)
where $E_{\bm{k},j}$ is the eigen energy of the BdG Hamiltonian. Two results are identical to each other when the transition to the superconducting states is a second-order. For first-order transitions, the phase boundary is obtained from the free-energy minima. The transition to the superconducting state becomes a first-order for $T<0.556T_{0}$ [9, 17] because odd-frequency pairing correlation decreases the coefficient of the fourth order term in the Ginzburg-Landau expansion and changes its sign to negative at low temperatures [18]. The arrow on the vertical axis indicate the Pauli limit $H_{\mathrm{P}}$ [14, 4]. A Zeeman field seriously suppresses spin-triplet superconductivity in the parallel configuration.
Secondly, we consider a configuration $\bm{\alpha}\perp\bm{d}$, $\bm{\alpha}\perp\bm{H}$, and $\bm{H}\perp\bm{d}$, where the three vectors are perpendicular to one another. The anomalous Green’s function near the transition temperature is supplied in Appendix (28). Here we show the results after carrying out the summation over $\bm{k}$ by shifting the wavenumber $\bm{k}\to\bm{k}\mp\bm{K}$ for $\xi_{\pm}$,
$\frac{1}{V_{\mathrm{vol}}}\sum_{\bm{k}}$ $\check{F}_{\perp}(\bm{k},\omega_{n})=-\frac{N_{0}\pi}{|\omega_{n}|(\omega_{n}^{2}+\alpha^{2}+\mu_{\mathrm{B}}^{2}H^{2})}$
$\times\left[(\omega_{n}^{2}+\mu_{\mathrm{B}}^{2}H^{2})\bm{d}\cdot\hat{\bm{\sigma}}+\omega_{n}(\bm{\alpha}\times\bm{d})\cdot\hat{\bm{\sigma}}\hat{\rho}_{z}\right$.
$\left.+i(\bm{\alpha}\times\bm{d})\cdot\mu_{B}\bm{H}\hat{\rho}_{z}\right]i\hat{\sigma}_{y}(i)\hat{\rho}_{y}.$ (9)
The first term is the pairing correlation linked to the pair potential. The odd-frequency pairing correlation at the second term implies that the SOI suppresses superconductivity [6]. The third term plays a key role in the ZFIS. This term stabilizes the superconducting state at high Zeeman field because it belongs to even-frequency symmetry class. The linearized gap equation results in
$\ln\frac{T}{T_{0}}$ $+2\pi T\sum_{\omega_{n}>0}\frac{1}{\omega_{n}}\left[1-\frac{\omega_{n}^{2}+\mu_{\mathrm{B}}^{2}H^{2}}{\omega_{n}^{2}+\mu_{\mathrm{B}}^{2}H^{2}+\alpha^{2}}\right]=0.$ (10)
Comparing the results with Eq. (7) suggests that the large SOI $\alpha>\mu_{\mathrm{B}}H_{\mathrm{P}}$ deletes superconductivity at $\bm{H}=0$. In the HTH-T phase diagram in Fig. 1(b), superconductivity is absent at $\bm{H}=0$ because we choose such large SOI as $\alpha=1.5T_{0}$ and $2.5T_{0}$. Equation (10) also indicates that a Zeeman field screens such negative effects due to the SOI. Fig. 1(b) shows that the superconducting phase appears in large Zeeman fields and that the critical temperature increases with increasing Zeeman fields. The induced spin-singlet Cooper pairs at the last term in Eq. (9) stabilize the spin-triplet superconducting states at high Zeeman fields. We conclude that the interplay between the SOI and the Zeeman field realizes ZFIS. Eq. (10) suggests that $T_{c}$ goes to the $T_{0}$ for $\mu_{\mathrm{B}}H\gg\alpha$. However, our phenomenological theory cannot predict such limiting behavior. We have assumed in high magnetic fields that the attractive interaction between two electrons remains unchanged and that other conduction bands do not come to the Fermi level. In Table 1, we list the pairing correlations appearing in a SC and their pairing symmetry.
Next, we consider mixed situations between the parallel and the perpendicular configurations,
$\bm{\alpha}=$ $\alpha_{x}\bm{e}_{x}+\alpha_{y}\bm{e}_{y},$ (11)
$\bm{d}=$ $d_{x}\bm{e}_{x}+d_{z}\bm{e}_{z},$ (12)
and $\bm{H}=H\bm{e}_{x}$, where both the SOI and the pair potential have two components. The anomalous Green’s function has complicated structure as shown in Eq. (19). The three induced pairing correlations in Table 1 characterize the stability and the instability of the mixed superconducting state. The phase diagram is shown in Fig. 1(c), where we choose $\bm{g}=(g,0,g)$ in Eq. (24), fix $\alpha_{x}$ at 0.75 $T_{0}$, and determine the phase boundary from the minima of the free-energy. At $H=0$, odd-frequency spin-triplet Cooper pairing correlation proportional to $\bm{\alpha}\times\bm{d}\cdot\hat{\bm{\sigma}}$ make the superconducting states unstable. Both $\alpha_{x}$ and $\alpha_{y}$ suppress the pair potential $d_{z}$, whereas only $\alpha_{y}$ suppresses $d_{x}$. As a result, single-component superconducting states with $\bm{d}=(d_{x},0,0)$ are realized for small $H$. With increasing Zeeman fields from zero, the transition temperature first decreases because odd-frequency spin-singlet pairing correlation proportional to $\bm{d}\cdot\bm{H}$ suppresses such superconducting state. The results for $\alpha_{y}=1.25T_{0}$ show that superconductivity vanishes for $0.55T_{0}<\mu_{\mathrm{B}}H<1.12T_{0}$ even at $T=0$. However, another single component superconductivity with $\bm{d}=(0,0,d_{z})$ appears for high Zeeman fields $1.12T_{0}<\mu_{\mathrm{B}}H$. A Zeeman field does not affect $d_{z}$ component because of $\bm{d}\cdot\bm{H}=0$. In addition, the spin-singlet pairing correlation proportional to $\bm{\alpha}\times\bm{d}\cdot\bm{H}$ stabilize superconducting states because it increases the superfluid density. We conclude that spin-singlet Cooper pairs induced by the interplay between the SOI and a Zeeman field cause the reentrant superconductivity. A spin-singlet Cooper pair in high Zeeman fields is an object which does not align with institution. However, our calculation clearly indicates its existence. A recent experiment on UTe2 shows a sign of spin-singlet pairs in the high field phase [13].
Figure 2: The critical magnetic field $H_{c}$ is plotted as a function of temperature for a single-band spin-triplet superconductors with Rashba SOI. The pair potential is chosen as Eq. (16a) in (a) and Eq. (16b) in (b).
Finally, we demonstrate the reentrant superconductivity in a usual single-band SC in two-dimension,
$\check{H}_{\mathrm{N}}(\bm{k})=$ $\left(\frac{\bm{k}^{2}}{2m}-\mu\right)\hat{\sigma}_{0}+\bm{\alpha}_{\bm{k}}\cdot\hat{\bm{\sigma}}+\mu_{\mathrm{B}}{H}\hat{\sigma}_{x},$ (13)
$\hat{\Delta}(\bm{k})=$ $i\bm{d}_{\bm{k}}\cdot\hat{\bm{\sigma}}\,i\hat{\sigma}_{y},$ (14)
$\bm{\alpha}_{\bm{k}}=$ $\alpha\left(\bar{k}_{y}\bm{e}_{x}-\bar{k}_{x}\bm{e}_{y}\right),$ (15)
where we consider the Rashba SOI and $\bar{k}_{j}=k_{j}/k_{F}$ for $j=x,y$. We consider the two pair potentials as
$\bm{d}_{\bm{k}}=$ $d_{x}\bar{k}_{x}\bm{e}_{x}+d_{z}\bar{k}_{y}\bm{e}_{z},$ (16a)
$\bm{d}_{\bm{k}}=$ $d_{x}\bar{k}_{y}\bm{e}_{x}+d_{z}\bar{k}_{x}\bm{e}_{z}.\$(16b)
We assume that the attractive interaction at $p$-wave channel in the gap equation. The phase diagrams for Eqs. (16a) and (16b) are shown in Fig. 2(a) and (b), respectively. The results for $\alpha=1.5T_{0}$ in (a) show ususal single superconducting phase, whereas those in (b) show two separated superconducting phases. The phase diagram depends sensitively on the relative $\bm{k}$ dependence between $\bm{\alpha}_{\bm{k}}$ and $\bm{d}_{\bm{k}}$ for large $\alpha$. Such tendency seems to be weaker for smaller $\alpha$. The reentrant superconductivity can be seen in the results for $\alpha=1.2T_{0}$ and $\alpha=T_{0}$ in both Figs. 2(a) and (b).
Conclusion: We have studied the stability of a spin-triplet superconducting state described by a $\bm{d}$ vector in the presence of a spin-orbit interaction $\bm{\alpha}$ and a Zeeman field $\bm{H}$. The spin-orbit interaction for $\bm{\alpha}\times\bm{d}\neq 0$ and Zeeman fields for $\bm{H}\cdot\bm{d}\neq 0$ make the superconducting state unstable and decrease the transition temperature. The instability is characterized well by the appearance of odd-frequency Cooper pairs because they decrease the superfluid density. We demonstrate that a spin-triplet superconducting state exhibits the Zeeman-field-induced superconductivity when a scaler chiral product of $\bm{\alpha}\times\bm{d}\cdot\bm{H}$ remains finite. Under the condition, even-frequency spin-singlet even-parity Cooper pairs stabilize the spin-triplet superconductivity at high Zeeman fields. Based on the obtained results, we propose a Hamiltonian of superconducting states which indicate the reentrant superconductivity in Zeeman fields. Although the Jaccarino-Peter compensation has been a reasonable story for the Zeeman-field-induced superconductivity, our conclusion provides an alternative physical picture for the phenomenon. In addition, our theory can be applied also to superconductors preserving time-reversal symmetry in the absence of a magnetic field.
Acknowledgments: T. S. is supported by JST SPRING, Grant Number JPMJSP2119. S. I. is supported by a Grant-in-Aid for Early-Career Scientists (JSPS KAKENHI Grant No. JP24K17010).
Appendix A Solusion of Gor’kov equation
The BdG Hamiltonian is represented by
$H_{\mathrm{BdG}}(\bm{k})$ $=\begin{bmatrix}\xi_{-}+\bm{\alpha}\cdot\hat{\bm{\sigma}}+\bm{B}\cdot\hat{\bm{\sigma}}&0&0&i\bm{d}\cdot\hat{\bm{\sigma}}\,\hat{\sigma}_{y}\\ 0&\xi_{+}-\bm{\alpha}\cdot\hat{\bm{\sigma}}+\bm{B}\cdot\hat{\bm{\sigma}}&-i\bm{d}\cdot\hat{\bm{\sigma}}\,\hat{\sigma}_{y}&0\\ 0&-i\bm{d}\cdot\hat{\bm{\sigma}}^{\ast}\,\hat{\sigma}_{y}&-\xi_{+}-\bm{\alpha}\cdot\hat{\bm{\sigma}}^{*}-\bm{B}\cdot\hat{\bm{\sigma}}^{*}&0\\ i\bm{d}\cdot\hat{\bm{\sigma}}^{\ast}\,\hat{\sigma}_{y}\ &0&0&-\xi_{-}+\bm{\alpha}\cdot\hat{\bm{\sigma}}^{*}-\bm{B}\cdot\hat{\bm{\sigma}}^{*}\end{bmatrix}$, (17)
$\xi_{\pm}$ $=\frac{1}{2m}(\bm{k}\pm\bm{K})^{2}-\mu,\quad\bm{B}=\mu_{\mathrm{B}}\bm{H},$ (18)
where $8\times 8$ matrix structure is derived from spin, valley and particle-hole degree of freedom of a quasiparticle. The Hamiltonian can be block-diagonalized into two $4\times 4$ Hamiltonians. The anomalous Green’s function is calculated as a solution of the Gor’kov equation
$\check{F}(\bm{k},\omega_{n})=$ $-\bm{d}^{2}\,\check{Z}^{-1}\left[(\omega_{n}^{2}+\hat{\xi}_{\bm{k}}^{2}+\bm{d}^{2}-\alpha^{2}+\bm{B}^{2})\,\bm{d}\cdot\hat{\bm{\sigma}}-2(\bm{B}\cdot\bm{d})\bm{B}\cdot\hat{\bm{\sigma}}+2(\bm{\alpha}\cdot\bm{d})\bm{\alpha}\cdot\hat{\bm{\sigma}}\right$.
$-2\hat{\xi}_{\bm{k}}\,\bm{\alpha}\cdot\bm{d}\hat{\rho}_{z}+2\,i\,\hat{\xi}_{\bm{k}}\,(\bm{B}\times\bm{d})\cdot\hat{\bm{\sigma}}$
$\left.+2i\,\omega_{n}\,\bm{B}\cdot\bm{d}+2\,\omega_{n}\,(\bm{\alpha}\times\bm{d})\cdot\hat{\bm{\sigma}}\,\hat{\rho}_{z}-2i\,\bm{\alpha}\times\bm{B}\cdot\,\bm{d}\,\hat{\rho}_{z}\right]\,i\hat{\sigma}_{y}\,(i\hat{\rho}_{y}),$ (19)
$\check{Z}(\bm{k},\omega_{n})=$ $\left\{\omega_{n}^{2}+\hat{\xi}_{\bm{k}}^{2}+\bm{d}^{2}-\bm{\alpha}^{2}+\bm{B}^{2}\right\}^{2}\bm{d}^{2}-4(\hat{\xi}_{\bm{k}}\bm{B}-i\omega_{n}\bm{\alpha}\hat{\rho}_{z})^{2}\bm{d}^{2}+4(\bm{\alpha}\cdot\bm{d})^{2}(\bm{B}^{2}+\bm{d}^{2})+4(\bm{B}\cdot\bm{d})^{2}(\bm{\alpha}^{2}-\bm{d}^{2})$
$+4(\bm{\alpha}\times\bm{B}\cdot\bm{d})^{2}-8(\bm{\alpha}\cdot\bm{d})(\bm{B}\cdot\bm{d})(\bm{\alpha}\cdot\bm{B}),$ (20)
$=$ $\hat{\rho}_{x}\,\check{Z}(-\bm{k},-\omega_{n})\,\hat{\rho}_{x}.$ (21)
The anomalous Green’s function is antisymmetric under the operation of interchanging two electron,
$\check{F}(\bm{k},\omega_{n})=-\hat{\rho}_{x}\,\check{F}^{\mathrm{T}}(-\bm{k},-\omega_{n})\,\hat{\rho}_{x},$ (22)
where $\mathrm{T}$ represents the transpose of the Pauli matrices for spin meaning the exchange of two spins and $\hat{\rho}_{x}\cdots\hat{\rho}_{x}$ represents the exchange of the two valleys. Eq. (21) indicates that $\check{Z}$ is symmetric under such operation. The three components on the first line in Eq. (19) represent spin-triplet odd-valley-parity symmetry class and are linked to the pair potential through the gap equation,
$i\left(\bm{d}\cdot\hat{\bm{\sigma}}\,\hat{\sigma}_{y}\right)_{\alpha,\beta}=$ $-T\sum_{\omega_{n}}\frac{1}{V_{\mathrm{vol}}}\sum_{\bm{k}}\sum_{\gamma,\delta}g_{\alpha\,\beta;\gamma\,\delta}\hat{F}^{(y)}_{\gamma,\delta}(\bm{k},\omega_{n}),$ (23)
$g_{\alpha\,\beta;\gamma\,\delta}=$ $\sum_{\nu=x,y,z}g_{\nu}\left(i\hat{\sigma}_{\nu}\,\hat{\sigma}_{y}\right)_{\alpha,\beta}\;\left(i\hat{\sigma}_{\nu}\,\hat{\sigma}_{y}\right)^{\ast}_{\gamma,\delta},$ (24)
where $\hat{F}^{(y)}$ is the $\hat{\rho}_{y}$ component of the anomalous Green’s function and $\bm{g}=(g_{x},g_{y},g_{z})$ represents the attractive interaction between two electrons at an $s$-wave channel. To draw Fig. 2, we assume an attractive interaction at $p$-wave channel as usual
$g_{\nu}(\bm{k}-\bm{k}^{\prime})=g_{\nu}\cos(\theta-\theta^{\prime}),\;\cos\theta=k_{x}/k_{F},\;\sin\theta=k_{y}/k_{F}.$ (25)
To our knowledge, the two components proportional to $\hat{\xi}_{\bm{k}}$ at the second line in Eq. (19) do not play any important role in stabilizing superconducting states. Namely, the presence of these components does not change $T_{c}$ at all. The first two components at the last line belong to the odd-frequency symmetry class and make the superconducting state unstable. [3]. A Zeeman field parallel to $\bm{d}$ and a SOI perpendicular to $\bm{d}$ generate such odd-frequency pairs from the pair potential. As a result, these components decrease $T_{c}$. The last component at the third line in Eq. (19) belongs to even-frequency spin-singlet even-valley-parity symmetry class. This component stabilizes a superconducting state at high Zeeman field. However, it appears only when the three vectors have a finite spin chiral product $\bm{\alpha}\times\bm{H}\cdot\,\bm{d}$.
The anomalous Green’s function for $\bm{\alpha}\parallel\bm{H}\parallel\bm{d}$ near the transition temperature is represented as
$\check{F}_{\parallel}(\bm{k},\omega_{n})=$ $-\frac{1}{2}\left[(\hat{X}_{+}^{-1}+\hat{X}_{-}^{-1})\bm{d}\cdot\hat{\bm{\sigma}}-(\hat{X}_{+}^{-1}-\hat{X}_{-}^{-1})d\right]i\hat{\sigma}_{y}\,(-i)\hat{\rho}_{y},$ (26)
$\hat{X}_{\pm}$= $(\omega_{n}\pm i\mu_{\mathrm{B}}H)^{2}+(\hat{\xi}_{\bm{k}}\mp\alpha\,\hat{\rho}_{z})+d^{2}.$ (27)
It is possible to analyze the symmetry of Cooper pairs even after summation over $\bm{k}$ because all Cooper pairs belong to even-parity $s$-wave symmetry class. The first term represents the pairing correlation belonging to even-frequency spin-triplet odd-valley parity class and is linked to the pair potential through the gap equation. A Zeeman field induces the pairing correlation belonging to odd-frequency spin-singlet odd-valley parity class as shown in the second term. The anomalous Green’s function for $\bm{d}\perp\bm{\alpha}$, $\bm{d}\perp\bm{H}$, and $\bm{H}\perp\bm{\alpha}$ near $T_{c}$ is calculated to be
$\check{F}_{\perp}(\bm{k},\omega_{n})=$ $-\check{Z}_{\perp}^{-1}\left[(\omega_{n}^{2}+\hat{\xi}_{\bm{k}}^{2}-\alpha^{2}+\mu_{\mathrm{B}}^{2}H^{2})\bm{d}\cdot\hat{\bm{\sigma}}-2i\hat{\xi}_{\bm{k}}\,\mu_{\mathrm{B}}\,\bm{d}\times\bm{H}\cdot\hat{\bm{\sigma}}\right$.
$\left.+2\omega_{n}\bm{\alpha}\times\bm{d}\cdot\hat{\bm{\sigma}}\hat{\rho}_{z}+2i\mu_{\mathrm{B}}\bm{\alpha}\times\bm{d}\cdot\bm{H}\hat{\rho}_{z}\right]i\hat{\sigma}_{y}(-i)\hat{\rho}_{y},$ (28)
$\check{Z}_{\perp}=$ $\hat{\xi}_{\bm{k}}^{4}+2\hat{\xi}_{\bm{k}}^{2}(\omega_{n}^{2}-\alpha^{2}-\mu_{\mathrm{B}}^{2}H^{2})+(\omega_{n}^{2}+\alpha^{2}+\mu_{\mathrm{B}}^{2}H^{2})^{2}.$ (29)
The SOI makes the superconducting state unstable because it generate odd-frequency pairing correlation at the second line in Eq. (28). The last term represents even-frequency spin-singlet even-valley parity Cooper pairs which stabilize the superconducting state at high Zeeman fields.